10. Let f(x1, x2, x3 6 22r3 Find stationary points and classify them as minimums or maximums ) xf+2%+3ag+ x1x22r123 0112r2
Answer #1

Sol1 2 2 foau)=^+2: 222- 22 22 10 Statinm piuk find Jit to Q are f. Paltial delivatives 2, 2 213 12 ^2 - 22 = o 2, A. 63 bo veiar find Vlue put the NOL, thaw Values neve are vdues Eijin the if MAxi huuM point positive ten mi ni mum if n the eisen Vlu2. 2-a 4-a 1 2 (2-)(-)(6)-)-6)-)- (2-) 24-10, +-(4-a) 2(2-1) = (e- {F. (1A +23) -a(a-o +23) -4+ +4al4 = ニ 0 24 10ta 2 10 23a